OpenStudy (anonymous):

Find the volume of a sphere with diameter 22 meters. A. 6272.19 m3 B. 5575.28m3 C. 2027.37 m3 D. 44602.24 m3

OpenStudy (mayankdevnani):

do you know the volume of sphere?

OpenStudy (mayankdevnani):

do you know @alainabbyboo22

OpenStudy (anonymous):

no i don't .

OpenStudy (mayankdevnani):

(4/3) r^3=volume

OpenStudy (mayankdevnani):

radius=diameter/2

OpenStudy (mayankdevnani):

so r=11m

OpenStudy (mayankdevnani):

\[\frac{4}{3}\times \frac{22}{7}\times11^3\]

OpenStudy (mayankdevnani):

can you solve it @alainabbyboo22

OpenStudy (anonymous):

B. 5575.28m3 is my answer !